tyson:

Solve for x: 7x + 15 = 38

JSVSL7:

1. Use subtraction property of equality 2. Use division property of equality boom isolated x

AngeI:

This was already answered c:

tyson:

it wasnt could u tell me again

AngeI:

It was thou

tyson:

well i had to go when u gave me the anwser and i never got it so could u tell me agagain

AngeI:

It's still on the previous post that you said "Im back" on

tyson:

oh ok

russkaya05:

Simplifying 7x + 15 = 38 Reorder the terms: 15 + 7x = 38 Solving 15 + 7x = 38 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-15' to each side of the equation. 15 + -15 + 7x = 38 + -15 Combine like terms: 15 + -15 = 0 0 + 7x = 38 + -15 7x = 38 + -15 Combine like terms: 38 + -15 = 23 7x = 23 i think thats all hoo finally

AngeI:

Then you divide both sides by 7 to isolate x